Find an equation for the line through the point (7−9) with x intercept 3.
Answer by stanbon(75887)

Find an equation for the line through the point (7−9) with x intercept 3.
------
x-intercept: (3,0)
point: (7,-9)
---
slope = (-9-0)/(7-3) = -9/4
----
Form: y = mx + b
Solve for "b":
0 = (-9/4)3 + b
b = 27/4
-----
Equation:
y = (-9/4)x + 27/4
===============================
Cheers,
Stan H.
======================